A 30 nC charge experiences a 0.044 N electric force.A 30 nC charge experiences a 0.044 N electric force. Part A What is the magnitude of electric field at the position of this charge? Express your answer with the appropriate units. E- 1 Value Submit Previous Answers Request Answer X Incorrect; Try Again; 2 attempts remaining Enter your answer using units of electric field.
Answer #1

Relation Between Electric field and electric force is given by:

F = |q|*|E|

|E| = Magnitude of electric field = F/|q|

Given that

F = 0.044 N

|q| = 30 nC = 30*10^-9 C

So,

|E| = F/|q|

|E| = 0.044/(30*10^-9)

|E| = 1466666.67 N/C

In two significant figures

|E| = 1.5*106 N/C = 1.5 MN/C

Where 1*106 = 1 M

Use Correct units in your final answer (N/C)